Solve 36x2 -12ax + (a2 – b2) = 0
Answer by stanbon(75887)

Solve 36x2 -12ax + (a2 – b2) = 0
------------
[36x2 -12ax + a2] – b2 = 0
----
(6x-a)^2 - b^2 = 0
-----
Factor:
[6x-a-b][6x-a+b] = 0
----
x = (a+b)/6 or x = (a-b)/6
------------------------------
Cheers,
Stan H.